Fermi National Laboratory
 

Accelerator Update for October 25, 2007 - October 26, 2007

 

Thursday October 25

The day shift began with Operations monitoring an Antiproton Source (Pbar) stack (12.2E10) and a Recycler stash (6.8E10), with the Tevatron (TeV), Main Injector (MI) and Recycler conducting studies, and with beam to the SciBooNE / MiniBooNE experiments.

At 8:50 AM, Recycler experts reported a problem with the stochastic cooling. An expert investigated.

Operations reported at 10:06 AM that FESS technicians successfully switched the Pbar 95 LCW pumps.

A MI expert began a two-hour coalescing study at 10:11 AM.

Beam to Sci/MiniBooNE tripped off at 11:32 AM due to a change of beam positions at VP860. Operations turned off the auto tune program and put the beam position back to its correct position. An expert will investigate the problem later.

An instrumentation expert reported that he had completed his work on a MI DCCT to reduce its noise problem at 12:14 PM.

The Neutron Therapy Facility (NTF) reported at 1:17 PM that they had completed treating their last patient.

At 1:28 PM, MI RF experts reported that they were fixing an LCW leak on MIRF13.

Operations reported at 2:19 PM that they had prepared the Meson M01-M05 enclosures for access.

Operations reported at 3:35 PM that they had search and secure the MuCool (MTA) enclosure.

TeV experts began a study at 4:42 PM.

Pbar experts resumed stacking at 9:09 PM.

At 9:39 PM, Operations reported a MI vacuum burst. The vacuum recovered and Pbar resumed stacking.

Pbar experts turned over stacking to the Operations group at 11:51 PM.

Friday October 26

The midnight shift began with Pbar stacking (33.2E10), with Operations monitoring the stash (3.5E10), with TeV and MI experts conducting studies, and with Sci /MiniBooNE taking beam.

At 1:03 AM, the 750 keV area buncher tripped off for the first of two times in four hours. It reset both times.

At 2:36 AM, LRF5 tripped off. It reset.

At 2:45 AM, a Linac Klystron RF station (KRF1) tripped off on a LCW flow indication. The station went out of phase, but eventually locked back in.

At 3:10 AM, a Pbar Debuncher quadrupole power supply (D:Q729) tripped off due to a phase imbalance and would not reset. Operations called in an expert.

A Pbar expert reported at 4:06 AM that he had replaced the phase imbalance relay for Q729. The supply was turned back on and stacking resumed.

At 5:53 AM, TRF1 tripped off and would not reset and its heaters would not turn on. Experts said that the RF station being off was not interfering with their studies. Operations informed an expert.

At 7:33 AM, the 750 KeV area buncher tripped off again. Experts investigated and reset the device.

The Plans for Friday and the Weekend

The current plan is for the accelerators to continue with studies and commissioning.

Machine Status

Linac: They will need 30 minutes to repair the 750 KeV buncher trips.

Booster: They were stable all night, and then the beam began to suffer a ramp problem, which they are investigating.

Main Injector: They reported that the auto tune program, tuning beam to Sci/MiniBooNE, was suffering from a timing problem; they have temporarily corrected this problem. The vacuum bursts they had, caused vacuum pumps to trip off. They will need an access to switch these pumps back on. They would also like to conduct aperture scans today.

Antiproton Source: They reported that their stacking efficiency was good. They would like 4-8 hours to conduct an Accumulatorr study.

Recycler: The Pelletron experts report that they are still having controls problems. They would like some more antiprotons to tune Electron Cooling.

Tevatron: Experts report that their low beta collision points are good, but they still need to work on the ramp.

Switchyard 120: Experts reported that they ran their power supplies and had their safety walk through yesterday. They would like to try and get beam to the MT6 enclosure today. They don’t want any beam this weekend. Next week they need access to the F3 Manhole to mark the positions for two quadrupole magnets.

Sci/MiniBooNE: They have been getting good beam. MiniBooNE is off due to problems with their air conditioning system (HVAC).

NuMI: They are conditioning their new chiller system. Their dehumidifier has been installed and they are pulling control wires. They are on schedule to take beam on November 12 th.

CDF: They would like 2-3 hours of no beam so they can calibrate their silicon system.

D0: They are presently accessing their hall to work on a silicon sequencer power supply and other small jobs.

FESS: They are preparing to conduct the required yearly Wilson Hall fire pump test next week. This test will take 3-4 hours.

Controls: Experts are working on the Recycler controls scope problems and with scope problems elsewhere. They would like to work on TeV block transfers, but the beam must be off for this work.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/26/2007    email Fermilab